In how many ways can 3 men be selected out of 15 men if 2 of the men are to be excluded from every selection?
1
Expert's answer
2020-03-06T10:24:00-0500

It is given that 2 of 15 men are to be excluded from every selection.

Therefore, we must select 3 men out of 15-2=13 men.

We can do it in "_{13} C_3 =\frac{13\times 12\times 11}{3\times 2\times 1}=286" ways.

Answer: "286" ways.